# Broken Expectations: The Effects of Modelling Assumptions on the Inferred Dark Matter Distribution in the Milky Way’s Satellites

### v1

@@ 3.2 Results with pyGravSphere
The recovered density profile for Sculptor is somewhat reminiscent of the results for Carina, also with a preference for more radial velocity anisotropy (see Figure B2). We note that Fornax, Sculptor and Carina also have the largest kinematic samples of the dwarfs we consider, so it is expected that the biases inherent in applying the spherical Jeans equation to tidally perturbed systems would become more prominent (Read et al., 2021). Moreover, it is clear from Figure 7 that Fornax, Sculptor and Carina are the three galaxies in our sample that still exhibit a Plummer-characteristic rise of the velocity dispersion in the outer regions (see Section A of the Appendix), while Draco and Ursa Minor are nearly flat or declining. We suspect that this rise in the velocity dispersion profile is, in part, what leads pyGravSphere to infer higher densities near the half-light radius, which relates to the accuracy of inference of the central density. We discuss this more in Section C and D of the Appendix, where we also show that for rising 4-th velocity moments, the method of Read et al. (2018) tends to over- or underestimate the virial shape parameters depending on the spatial extent of the data.
We have additionally applied the Jeans analysis to the nearest pericentre and apocentre snapshots of the dwarfs. The comparison of the accuracy of the recovered profiles is shown in the subpanels of Figures 8 and 9. Overall, we see that the quality of the density profile recovery is consistent with that of the present-day snapshots. The main discrepancy is in the outskirts of the dwarfs, where the density inferred at apocentre is overestimated compared to the pericentre. This may seem counterintuitive, but the majority of the mass loss occurs after the pericentric passage is complete, such that the outer density slope is higher at the next apocentre than at pericentre (Peñarrubia et al., 2009). Since our priors assume density slopes between 0 and 3, the inferred density is, therefore, more inaccurate at apocentre (if the apocentre snapshot is after the pericentre). Note that this trend is reversed for Carina in the heavy potential, where the pericentre snapshot is after the apocentre one.
@@
Our results may seem at odds with those of Genina et al. (2020), who tested pyGravSphere on Fornax-like systems in the APOSTLE cosmological simulations and generally found a minor underestimation of the density with pyGravSphere at all radii; however, the systems that they consider depart from sphericity, which drives overestimation or underestimation of the enclosed mass, have smaller samples of stellar particles (, on par with our models of Draco and Ursa Minor) and lack the spatial resolution to analyse densities recovered below pc.
@@ 3.3 Dark matter central densities
As we have shown in Fig. 1, some of our models in the initial conditions are outliers in the mass-concentration relation, in particular the two models for Draco and the ‘heavy’ potential models for Ursa Minor and Sculptor. Their initial inner densities are shown with blue arrows in Figure 10. Through mass loss due to tidal stripping, their present-day inner densities are in good agreement with the mass-concentration relation, however. In the future, with more realistic models of the Milky Way potential and the LMC, it would be informative to explore whether NFW haloes are compatible with the present-day densities of Milky Way’s satellites when tidal stripping is taken into account. Our results suggest that insignificant tidal stripping (i.e. a ‘light’ Milky Way potential) can put the observed kinematics of Milky Way’s dwarfs in agreement with CDM, but would still require Draco and Fornax to be outliers in the mass-concentration relation (assuming 0.1 dex scatter in the relation at the low-mass end). It is also unclear how this result fits into the ‘tidal stirring’ scenario of dwarf spheroidal formation from initially disky galaxies (Mayer et al., 2001), as it seems unlikely that Milky Way’s tides can affect the stellar populations of the five dwarfs considered in this work significantly (Di Cintio et al., 2024), especially if the infall time of these dwarfs is late (Battaglia et al., 2022). On the other hand, the preferred stripping of stars on prograde tangentially anisotropic orbits may still reconcile this apparent inconsistency (Read et al., 2006; Łokas et al., 2015), but will require further tests with simulations to confirm this. Sudden loss of gas due to ram pressure and subsequent tidal effects may also alleviate the tension (Hammer et al., 2024).
We additionally note that, in part, the agreement we see here results from pyGravSphere’s tendency to underestimate the central densities of cuspy haloes hosting a Plummer stellar distribution. This is particularly apparent with Fornax in the ‘heavy’ and ‘light’ potential, with an additional reduction of the central density due to tides (Genina et al., 2022). We have explored modelling our simulated systems with the Zhao (1997) dark matter density model available in pyGravSphere and found a more faithful recovery of the inner density profile (purple and pink points in Fig. 10). This is partly because the Zhao (1997) parameterisation is less flexible in its ability to model small dark matter cores, thus forcing the density profile towards a more cuspy solution, which in our case is the correct one, but it may not be in nature. Despite this apparent success of the Zhao (1997) model in this case, we would argue for more non-parametric approaches to dynamical modelling, which will more likely return accurate solutions in the future as distances and proper motions become more available for individual stars in nearby dwarf spheroidals.
@@ 3.4 J-Factor Results
We calculate the J-factors of a sample of MCMC chain steps for each dSph directly from the Jeans analysis results, as well as the ‘true’ J-factors of the corresponding snapshots. Figure 11 summarises the results for the dSphs at present-day snapshots and when the dSphs were closest to their peri- and apocentres in the ‘light’ and ‘heavy’ Milky Way potential.
As a general trend, we find that the J-factors are typically underestimated for our simulated systems, which we attribute to the lower estimated central densities. Exceptions to this are Carina and Sculptor in the light potential, where there is a significant overestimation of the density within the central 1 kpc. The discrepancy in the outer profiles, which is primarily caused by our priors, would correspond to an insignificant contribution to the J-factors since regions below dominate the contribution (Springel et al., 2008b). At apocentre, the J-factors are likely to be underestimated in both the ‘light’ and ‘heavy’ potential. This is likely the result of the ‘stretching’ of the dwarf along the line of sight at apocentre that increases the ‘true’ J-factor. For present-day snapshots, the true value is typically contained within the 95 per cent confidence intervals. The difference between the inferred J-factors and the ‘true’ ones is generally small but not negligible. It is also clear that in some cases, the errors on the J-factors are underestimated. While our J-factors computed with pyGravSphere are in good agreement with those of Alvarez et al. (2020), who also used the GravSphere technique, our results suggest that the ‘true’ error on the J-factors should be at least of order , based on our modelling of tidally stripped and initially spherically-symmetric systems with pyGravSphere.
@@ 4 Conclusions
i) NFW haloes with Plummer stellar distributions are able to match the projected half-light radii and velocity dispersions within the half-light observed in Fornax, Carina, Sculptor, Draco and Ursa Minor at the present day. For the ‘light’ potential, the NFW haloes are consistent with the mass-concentration relation in CDM, however in the ‘heavy’ potential, Draco and Ursa Minor would need to be outliers from the mass-concentration relation. There is additionally a discrepancy for Fornax in the ‘heavy’ potential.
@@
iii) In the ‘light’ Milky Way potential, the part of the line-of-sight velocity profile that rises is generally maintained for our sample of dwarfs, while in the ‘heavy’ potential, the profile flattens or even steepens in the outskirts when only the bound stellar particles are considered, suggesting a possibility of inferring the mass of the Milky Way based on the internal kinematics of its satellites.
iv) When the Wolf et al. (2010) mass estimator is applied to these systems, the accuracy of the estimator depends on how the velocity-dispersion is derived. If the luminosity-weighted line-of-sight velocity dispersion is used, the accuracy of the estimator is affected by the rising or falling shape of the velocity dispersion profile in the ‘light’ and ‘heavy’ potential, respectively. The inferred masses are larger when the velocity dispersion profile is rising and lower when they fall. In some cases, the accuracy of the estimator can be improved by instead using the velocity dispersion within the half-light radius in the mass estimator. Nevertheless, we found both forms of the estimator to be accurate to better than 10 per cent for our simulated dwarf spheroidals at the present day, though we emphasise that we consider initially spherical systems.
@@
vi) We applied the non-parametric Jeans code pyGravSphere (Read & Steger, 2017; Genina et al., 2020) to our sample of simulated dwarfs and found that the severity of tidal effects had no significant effect on the quality of the density profile recovery, aside from the ability to model the outer parts of the dark matter profile, which is partially related to the density slope priors that we set. Only for Ursa Minor in the ‘heavy’ potential, we have observed an inflated velocity dispersion at the peri- and apocentre snapshots, which led to an inference of a more cuspy central density profiles. For the rest of the dwarfs, the quality of the recovered density profiles has followed the pattern already known with similar, isolated, mocks with a typical underestimation of the density at small radii and overestimation at large radii (Read & Steger, 2017; Read et al., 2021). This is the result of the “smoothness” constraint on the broken power-law density profiles, combined with the tendency for the virial shape parameters to be either under or over-estimated if the true extent of the maximum possible sample of observational data is not fully known (see the Appendix C and the discussion in Read et al. 2018). We suggest that the virial shape parameters estimated only for the available extent of observational data (Richardson & Fairbairn, 2014), without any extrapolation, should be used as lower limits on the virial shape parameters of pyGravSphere models.
@@
It would, of course, be interesting to confirm our findings with simulations that account for the influence of the Large Magellanic Cloud on dwarf spheroidal orbits as well as the Milky Way’s mass evolution (Vasiliev et al., 2021). Our models of the dwarf spheroidals are also not fully realistic, as they do not take into account the possible initial disky configuration of the dwarfs (Mayer et al., 2001), their present-day asphericity (Shchelkanova et al., 2021), rotation (Martínez-García et al., 2021), multiple kinematically-distinct stellar populations (Tolstoy et al., 2004; Battaglia et al., 2006; Pace et al., 2020) or the influence of gas and star formation on their potential wells (Pontzen & Governato, 2012). We have also not considered initially cored dark matter density distributions (Kazantzidis et al., 2013), and thus, it is unclear whether pyGravSphere would recover those accurately in the presence of tidal effects.
@@ Acknowledgements
KT and AG would like to thank Sherry Suyu and Volker Springel, without whom this work would not have been possible and Thorsten Naab for continued support. Additionally, the authors would like to thank Simon White and Klaus Dolag for helpful discussions and Justin Read for providing useful feedback on the manuscript and explaining the effect of the “smoothness” constraint on the broken power-law model. This research was carried out using the High-Performance Computing resources of the FREYA cluster at the Max Planck Computing and Data Facility (MPCDF, https://www.mpcdf.mpg.de) in Garching, operated by the Max Planck Society (MPG).
@@ Appendix A Line-of-sight velocity dispersions of Plummer profiles in NFW haloes
In Figure A1, we show the shape of the line-of-sight velocity dispersion profiles of Plummer stellar distributions residing in NFW haloes, assuming full isotropy. The profiles were obtained by solving the spherical Jeans equation. We note that a fully isotropic solution is not permitted by the condition of the non-negativity of the distribution function (Sánchez Almeida et al., 2023); however, since we allow our generated models to settle in equilibrium, they acquire some anisotropy (see Figures B1 and B2). We found that radial anisotropy in the outskirts tends to lower the velocity dispersion of the outer ‘bump’, while radial anisotropy in the inner region increases the velocity dispersion at the centre. Tangential anisotropy in the outskirts causes the velocity dispersion in the outer regions to rise. The ‘bump’ beyond the half-light radius is preserved with the change in the anisotropy profile but can be wiped away by tidal forces (see Figure 3), or may not be fully captured depending on the extent of the observational data. Variations in the mass profile can, of course, also lead to changes in the shape of the profile, with the classic mass-anisotropy degeneracy.

### v2

@@ 3.2.1 Effects of modelling assumptions
The recovered density profile for Sculptor is somewhat reminiscent of the results for Carina, also with a preference for more radial velocity anisotropy (see Figure B2). We note that Fornax, Sculptor and Carina also have the largest kinematic samples of the dwarfs we consider, so it is expected that the biases inherent in the modelling would become more prominent (Read et al., 2021). Moreover, it is clear from Figure 7 that Fornax, Sculptor and Carina are the three galaxies in our sample that still exhibit a Plummer-characteristic rise of the velocity dispersion in the outer regions (see Section A of the Appendix), while Draco and Ursa Minor are nearly flat or declining. We suspect that this rise in the velocity dispersion profile, the peak of which approximates well the scale radius of the halo (see Appendix A), is, in part, what leads pyGravSphere to infer higher densities near the half-light radius, which lowers the inferred central density. We discuss this further in the Supplementary Materials, where we also show that for rising 4-th velocity moments, the method of Read et al. (2018) tends to over- or underestimate the virial shape parameters depending on the spatial extent of the data, which affects the constraint on the outer density profile and, through the parameterisation of the model, the inferred densities in the inner regions.
3.2.2 Effects of tides
We have additionally applied the Jeans analysis to the nearest pericentre and apocentre snapshots of the dwarfs. The comparison of the accuracy of the recovered profiles is shown in the subpanels of Figures 8 and 9. Overall, we see that the quality of the density profile recovery is consistent with that of the present-day snapshots. Our analysis with the Wolf et al. (2010) mass estimator, even in the extreme case of Ursa Minor, also predicts variation of below the 10 per cent level for our sample of dwarf galaxies. The main discrepancy is in the outskirts of the dwarfs, where the density inferred at apocentre is overestimated compared to the pericentre. This may seem counterintuitive, but the majority of the mass loss occurs after the pericentric passage is complete, such that the outer density slope is higher at the next apocentre than at pericentre (Peñarrubia et al., 2009). Since our priors assume density slopes between 0 and 3, the inferred density is, therefore, more inaccurate at apocentre (if the apocentre snapshot is after the pericentre). Note that this trend is reversed for Carina in the heavy potential, where the pericentre snapshot is after the apocentre one.
@@ 3.2.2 Effects of tides
Our results may seem at odds with those of Genina et al. (2020), who tested pyGravSphere on Fornax-like systems in the APOSTLE cosmological simulations and generally found a minor underestimation of the density with pyGravSphere at all radii; however, the systems that they consider depart from sphericity, which drives overestimation or underestimation of the enclosed mass, have smaller samples of stellar particles (, on par with our models of Draco and Ursa Minor), with the spatial extent of the data limited to within 2 half-light radii, and lack the spatial resolution to analyse densities recovered below pc. The overestimation of the density at large radii due to the dark matter density log-slopes being limited to 3 is in agreement with their results.
Previous works of Battaglia et al. (2015) and Iorio et al. (2019) are the most similar in nature to ours, considering Fornax and Sculptor dwarf galaxies, respectively, tidally stripped in the Milky Way potential and on orbits constrained by the measured proper motions. Both works additionally apply Jeans dynamical modelling to their simulated dwarfs at the present-day snapshots. While there are some differences in the modelling choices, in particular, the details of the Milky Way potential and the dark matter distribution of the dwarf galaxies, both of these previous studies agree with our conclusions that for Sculptor and Fornax the assumption of dynamical equilibrium still holds. The work of Iorio et al. (2019) in particular has explored the most eccentric orbits permitted by proper motion errors and have taken a more observational approach to selecting the kinematic sample in their simulations for the dynamical modelling (based on selection function for observational data as opposed to considering all bound particles), they still conclude that tidal effects are not significant on the inferred density and mass distributions. Similar conclusions were reached in Read et al. (2018), who applied the GravSphere model to mock data generated from simulations of the properties and orbit of Draco. They found that GravSphere was able to faithfully recover the density profile of Draco after 10 Gyr of evolution in the Milky Way potential. As such, the results presented here are in good agreement with previous work, despite differences in modelling choices.
The work of Muñoz et al. (2008) models the orbital evolution of Carina dwarf spheroidal and finds a significant impact of tides, although they do not apply dynamical modelling to their modelled satellites. The difference from our results likely stems from the fact that they model the dark matter distribution of Carina with particles and using a cored Plummer (1911) model. Both of these choices would result in enhanced physical and artificial tidal disruption of their modelled satellites (van den Bosch & Ogiya, 2018; Errani et al., 2023).
To conclude, we find that for the sample of 5 initially spherical dwarfs in 2 Milky Way potentials considered in this work, the assumption of dynamical equilibrium is a reasonable one; instead, the inaccuracies we found stem from the parameterization of the dynamical models. As we only considered bound tracers in our analysis, we suggest that the presence of unbound stars in kinematic samples pose a greater risk to the accuracy of mass distributions recovered with spherical Jeans models for these dwarfs than the perturbations to the dynamical equilibrium (Klimentowski et al., 2007; Łokas et al., 2010).
@@ 3.3 Dark matter central densities
As we have shown in Fig. 1, some of our models in the initial conditions are outliers in the mass-concentration relation, in particular the two models for Draco and the ‘heavy’ potential models for Ursa Minor and Sculptor. Their initial inner densities are shown with blue arrows in Figure 10. Through mass loss due to tidal stripping, their present-day inner densities appear in good agreement with the mass-concentration relation, however. In the future, with more realistic models of the Milky Way potential and the LMC, it would be informative to explore whether NFW haloes are compatible with the present-day densities of Milky Way’s satellites when tidal stripping is taken into account. Our results suggest that insignificant tidal stripping (i.e. a ‘light’ Milky Way potential) can put the observed kinematics of Milky Way’s dwarfs in agreement with CDM, but would still require Draco and Fornax to be outliers in the mass-concentration relation (assuming 0.1 dex scatter in the relation at the low-mass end). It is also unclear how this result fits into the ‘tidal stirring’ scenario of dwarf spheroidal formation from initially disky galaxies (Mayer et al., 2001), as it seems unlikely that Milky Way’s tides can affect the stellar populations of the five dwarfs considered in this work significantly (Di Cintio et al., 2024), especially if the infall time of these dwarfs is late (Battaglia et al., 2022). On the other hand, the preferred stripping of stars on prograde tangentially anisotropic orbits may still reconcile this apparent inconsistency (Read et al., 2006; Łokas et al., 2015), but will require further tests with simulations to confirm this. Sudden loss of gas due to ram pressure and subsequent tidal effects may also alleviate the tension (Hammer et al., 2024; Wang et al., 2024).
We additionally note that, in part, the agreement we see here results from pyGravSphere’s tendency to underestimate the central densities. This is particularly apparent with Fornax in the ‘heavy’ and ‘light’ potential, with an additional reduction of the central density due to tides (Genina et al., 2022). We have explored modelling our simulated systems with the Zhao (1997) dark matter density model available in pyGravSphere and found a more faithful recovery of the inner density profile (purple and pink points in Fig. 10). This is partly because the Zhao (1997) parameterisation is less flexible in its ability to model small dark matter cores or tidal truncation, thus forcing the density profile towards a more cuspy solution, which in our case is the correct one, but it may not be in nature. Despite this apparent success of the Zhao (1997) model in this case, we would argue for more non-parametric approaches to dynamical modelling, which will more likely return accurate solutions in the future as distances and proper motions become more available for individual stars in nearby dwarf spheroidals.
@@ 3.4 J-Factor Results
We calculate the J-factors of a sample of MCMC chain steps for each dSph directly from the Jeans analysis results, as well as the ‘true’ J-factors of the corresponding snapshots. The top panel of Figure 11 summarises the results for the dSphs at present-day snapshots and when the dSphs were closest to their peri- and apocentres in the ‘light’ and ‘heavy’ Milky Way potential.
As a general trend, we find that the J-factors are typically underestimated for our simulated systems, which we attribute to the lower estimated central densities. Exceptions to this are Carina and Sculptor in the light potential, where there is a significant overestimation of the density within the central 1 kpc. The discrepancy in the outer profiles, which is primarily caused by our priors, would correspond to an insignificant contribution to the J-factors since regions below dominate the contribution to the J-factor in NFW haloes (Springel et al., 2008b). At apocentre, the J-factors are likely to be underestimated in both the ‘light’ and ‘heavy’ potential. This is likely the result of the dwarfs orienting along the line of sight at apocentre that increases the ‘true’ J-factor. This is evident in the middle panel of Fig. 11, where we show the comparison between the ‘true’ J-factors calculated only for the bound particles and those computed using all particles within the solid angle. Although the difference is rather small due to significantly lower densities of the stripped dark matter, it is clear that in the ‘light’ potential, the stripped debris tends to align with the line of sight at apocentre. In the ‘heavy’ potential, this difference is less significant as the dark matter in the dwarfs has already undergone significant stripping, and they lose progressively smaller fractions of their mass.
For present-day snapshots, the true value is typically contained within the 95 per cent confidence intervals. The difference between the inferred J-factors and the ‘true’ ones is generally small but not negligible. It is also clear that in some cases, the errors on the J-factors are underestimated. While our J-factors computed with pyGravSphere are in good agreement with those of Alvarez et al. (2020) (bottom panel of Fig. 11), who also used the GravSphere technique, our results suggest that the ‘true’ error on the J-factors should be at least of order , based on our modelling of tidally stripped and initially spherically-symmetric systems with pyGravSphere.
@@ 4 Conclusions
i) NFW haloes with stellar distributions that are well-fit by a Plummer profile are able to match the projected half-light radii and velocity dispersions within the half-light observed in Fornax, Carina, Sculptor, Draco and Ursa Minor at the present day. For the ‘light’ potential, the NFW haloes are consistent with the mass-concentration relation in CDM, however in the ‘heavy’ potential, Draco and Ursa Minor would need to be outliers from the mass-concentration relation. There is additionally a discrepancy for Fornax in the ‘light’ potential.
@@
iii) In the ‘light’ Milky Way potential, the part of the line-of-sight velocity profile that rises is generally maintained for our sample of dwarfs, while in the ‘heavy’ potential, the profile flattens or even steepens in the outskirts when only the bound stellar particles are considered, suggesting a possibility of inferring the mass of the Milky Way based on the velocity dispersion of its satellites. This prospect is, however, complicated by the need to obtain kinematic data samples that are free from contamination by unbound stars.
iv) When the Wolf et al. (2010) mass estimator is applied to these systems, the accuracy of the estimator depends on how the velocity-dispersion is derived. If the luminosity-weighted line-of-sight velocity dispersion is used, the accuracy of the estimator is affected by the rising or falling shape of the velocity dispersion profile in the ‘light’ and ‘heavy’ potential, respectively. The inferred masses are larger when the velocity dispersion profile is rising and lower when they fall. In some cases, the accuracy of the estimator can be improved by instead using the velocity dispersion within the half-light radius in the mass estimator. Nevertheless, we found both forms of the estimator to be accurate to better than 10 per cent for our simulated dwarf spheroidals at the present day (see also Errani et al. 2018), though we emphasise that we consider initially spherical systems.
@@
vi) We applied the non-parametric Jeans code pyGravSphere (Read & Steger, 2017; Genina et al., 2020) to our sample of simulated dwarfs and found that the severity of tidal effects had no significant effect on the quality of the density profile recovery when analysing only bound stellar tracers, aside from the ability to model the outer parts of the dark matter profile, which is partially related to the density slope priors that we set. We conclude that, for the sample of dwarfs considered in this work, the presence of unbound stars in kinematic samples likely poses a greater challenge to Jeans models than the dynamical perturbations due to tides themselves.
We also found a typical underestimation of the central density (Read & Steger, 2017; Read et al., 2021) which arises due to the inability of the broken power-law model to fit the dark matter halo profile, combined with the “smoothness” constraint on the broken power-law density profiles. The tendency for the virial shape parameters to be either under- or over-estimated if the true extent of the galaxy is not fully known (see the Supplementary materials and the discussion in Read et al. 2018) can worsen this issue. We thus suggest that dynamical models with pyGravSphere should either keep the power-law break radii as free parameters of the model, or to increase the number of bins until the convergence in the obtained result is reached (though as Read & Steger (2017) point out, an increase in the number of bins should also be motivated by the quality and quantity of the data). Additionally, we suggest that the virial shape parameters should be estimated only for the available extent of observational data (Richardson & Fairbairn, 2014) and used as lower limits on the virial shape parameters of pyGravSphere models.
@@
ix) We have made the snapshots of simulations analysed in this work publicly available and hope that these can serve as mock datasets to test alternative dynamical modelling techniques in the future. It would be straightforward to modify these mocks to e.g. include a contaminating sample of background Milky Way stars, model stellar binaries and explore different observational sample selections.
@@
It would, of course, be interesting to confirm our findings with simulations that account for the influence of the Large Magellanic Cloud on dwarf spheroidal orbits as well as the Milky Way’s mass evolution (Buist & Helmi, 2014; Vasiliev et al., 2021). Our models of the dwarf spheroidals are also not fully realistic, as they do not take into account the possible initial disky configuration of the dwarfs (Mayer et al., 2001), their present-day asphericity (Shchelkanova et al., 2021), rotation (Martínez-García et al., 2021), multiple kinematically-distinct stellar populations (Tolstoy et al., 2004; Battaglia et al., 2006; Pace et al., 2020) or the influence of gas and star formation on their potential wells (Pontzen & Governato, 2012). We have also not considered initially cored dark matter density distributions (Kazantzidis et al., 2013), and thus, it is unclear whether pyGravSphere would recover those accurately in the presence of tidal effects.
@@ Acknowledgements
KT and AG would like to thank Sherry Suyu and Volker Springel, without whom this work would not have been possible and Thorsten Naab for continued support. Additionally, the authors would like to thank Simon White and Klaus Dolag for helpful discussions and Justin Read for providing useful feedback on the manuscript and explaining the effect of the “smoothness” constraint on the broken power-law model. We also thank the anonymous referee for insightful comments. This research was carried out using the High-Performance Computing resources of the FREYA cluster at the Max Planck Computing and Data Facility (MPCDF, https://www.mpcdf.mpg.de) in Garching, operated by the Max Planck Society (MPG).
@@ Appendix A Line-of-sight velocity dispersions of Plummer profiles in NFW haloes
In the top panel of Figure A1, we show the shape of the line-of-sight velocity dispersion profiles of Plummer stellar distributions residing in NFW haloes, assuming full isotropy. The profiles were obtained by solving the spherical Jeans equation. We note that a fully isotropic solution is not permitted by the condition of the non-negativity of the distribution function (Sánchez Almeida et al., 2023); however, since we allow our generated models to settle in equilibrium, they acquire some anisotropy (see Figures B1 and B2). The dotted lines show the half-light radii of our simulated dwarfs in the initial conditions, which align well with a ‘dip’ in the velocity dispersion profile. The dashed lines show the location of the NFW scale radius, , which slightly precedes the peak of the ‘bump’ in the velocity dispersion profile.
The middle panel shows that the radial anisotropy in the outskirts tends to lower the velocity dispersion of the outer ‘bump’, while radial anisotropy in the inner region increases the velocity dispersion in the centre. Tangential anisotropy in the outskirts causes the velocity dispersion in the outer regions to rise. The ‘bump’ beyond the half-light radius is preserved with the change in the anisotropy profile but can be wiped away by tidal forces, or may not be fully captured depending on the extent of the observational data. Variations in the mass profile can, of course, also lead to changes in the shape of the profile, with the classic mass-anisotropy degeneracy, as we demonstrate in the bottom panel for the case of cored and cuspy haloes, for which we change the steepness of the outer slope of the dark matter profile. In particular, outer slopes or 5, significantly reduce the amplitude of the outer ‘bump’. This likely explains why our ‘heavy’ models do not exhibit a strong ‘bump’ in the line-of-sight velocity dispersion, despite having tangential anisotropy in the outskirts (Read et al., 2006; Peñarrubia et al., 2009).